# Enhancing Sample Selection Against Label Noise by Cutting Mislabeled Easy Examples

### v2

@@ B.3 Model Architectures
B.4 Training Procedures and Hyperparameters
@@
To enhance the robustness of our sample selection model, we also incorporated certain strategies from prior works (Lin et al., 2024b; Li et al., 2024a), training two networks and each network learn from the other’s soft predictions and utilizing exponential moving averages to stabilize training. Weak data augmentation techniques were applied during training to improve generalization. These included random cropping with a padding of 4 pixels, random horizontal flipping, and normalization using the dataset-specific mean and standard deviation.
B.5 Sample Selection Mechanism
Building upon the Late Stopping strategy (Yuan et al., 2023), we iteratively select a confident subset $\mathcal{D}^{s}$ of training samples, progressively reducing mislabeled data and enhancing the model’s focus on clean samples. We identify early-learned samples based on their learning times. For each sample $(\mathbf{x}_{i},y_{i})$, we define its learning time $LT_{i}$ as the earliest epoch when the model’s prediction stabilizes:
$LT_{i}=\min\left\{T_{i}\ \big{|}\ \hat{y}_{i}^{T_{i}-2}=\hat{y}_{i}^{T_{i}-1}=\hat{y}_{i}^{T_{i}}=y_{i}\right\},$ (13)
where $\hat{y}_{i}^{t}$ denotes the model’s predicted label at epoch $t$.
@@ B.5 Sample Selection Mechanism
B.6 Algorithm
Algorithm 1 Iterative Sample Selection with Early Cutting
1:Training data $\mathcal{D}^{0}$; Number of iterations $I_{\text{rate}}$; Early cutting rate $\gamma$
2:Trained model parameters $\theta$
3:for $i=1$ to $I_{\text{rate}}$ do
4:     1. Base Sample Selection:
5:     for all samples $(\mathbf{x}_{j},y_{j})\in\mathcal{D}^{i-1}$ do
6:         Compute learning time $L_{j}$ using Eq. (13)
7:         Select subset $\mathcal{D}^{s}$ using learning time $L_{j}$
8:     end for
9:     2. Early Cutting:
10:     for all samples $(\mathbf{x}_{j},y_{j})\in\mathcal{D}^{s}$ do
11:         Select subset $\mathcal{D^{\prime}}^{s}$ with earliest $\frac{1}{\gamma}$ proportion of samples in $\mathcal{D}^{s}$ based on $L_{j}$
12:         Compute confidence score $c_{j}$ and gradient norm $g_{j}$
13:         Remove samples from $\mathcal{D^{\prime}}^{s}$ satisfying all the following conditions:
14:          (a) Large loss; (b) High confidence; (c) Low gradient norm.
15:         Update training set $\mathcal{D}^{i}$
16:     end for
17:end for
18:Final Training Phase: Train model on refined $\mathcal{D}^{I_{\text{rate}}}$ until convergence
19:return Trained model parameters $\theta^{*}$
B.7 Baselines and Competitors
We re-implemented these methods under the same experimental settings as our proposed method. When re-implementing CSGN (Lin et al., 2024b), we used the AdamW optimizer and a stepped decay learning rate schedule, as specified in the original code. Notably, CSGN (Lin et al., 2024b) cannot handle tasks with too many classes such as ImageNet-1k well.
B.8 Training Time and Computational Complexity
@@
Table 8: Comparison of total training hours and test accuracy (mean±$\pm$±std, CIFAR-100, ResNet-34, 40% Symmetric noise rate).
@@
B.9 Ablation Study on Each Component in Early Cutting Method
@@
B.10 Transferability of Default Parameters and Method Robustness
@@
CIFAR-10 Tiny-ViT Early Cutting (Ours) 75.42%

### v3

@@ D.3 Model Architectures
D.4 Training Procedures and Hyperparameters
@@
To enhance the robustness of our sample selection model, we also incorporated certain strategies from prior works (Lin et al., 2024b; Li et al., 2024), training two networks and each network learn from the other’s soft predictions and utilizing exponential moving averages to stabilize training. Weak data augmentation techniques were applied during training to improve generalization. These included random cropping with a padding of 4 pixels, random horizontal flipping, and normalization using the dataset-specific mean and standard deviation.
D.5 Sample Selection Mechanism
Building upon the Late Stopping strategy (Yuan et al., 2023), we iteratively select a confident subset $\mathcal{D}^{s}$ of training samples, progressively reducing mislabeled data and enhancing the model’s focus on clean samples. We identify early-learned samples based on their learning times. For each sample $(\mathbf{x}_{i},\tilde{y}_{i})$, we define its learning time $LT_{i}$ as the earliest epoch when the model’s prediction stabilizes:
$LT_{i}=\min\left\{T_{i}\ \big|\ \hat{y}_{i}^{E_{i}-2}=\hat{y}_{i}^{E_{i}-1}=\hat{y}_{i}^{E_{i}}=\tilde{y}_{i}\right\},$ (14)
where $\hat{y}_{i}^{t}$ denotes the model’s predicted label at epoch $e$.
@@ D.5 Sample Selection Mechanism
D.6 Baselines and Competitors
We re-implemented these methods under the same experimental settings as our proposed method. When re-implementing CSGN (Lin et al., 2024b) using only supervised learning for Table 2, 3, 4, and 5. We used the AdamW (Loshchilov and Hutter, 2019) optimizer and a stepped decay learning rate schedule, as specified in the original code. Notably, CSGN (Lin et al., 2024b) cannot handle tasks with too many classes such as ImageNet-1k well.
D.7 Training Time and Computational Complexity
@@
[…]
@@
D.8 Ablation Study on Each Component in Early Cutting Method
@@
D.9 Transferability of Default Parameters and Method Robustness
@@
CIFAR-10 TinyViT Early Cutting (Ours) 75.42%
@@
(a) Impact on model performance from mislabeled examples learned at different stages, using CIFAR-100.
Figure 6: Impact of mislabeled samples learned at different stages on model generalization performance.
(a) The speed at which pretrained models on CIFAR-100 learn mislabeled examples from different stages.
Figure 7: Comparison of how pretrained models learn mislabeled examples from different learning stages.
Appendix E Evolution of Feature Space and Distance Ratios
To supplement the analysis in the main text (Figure 4(a)), which shows the feature space and distance ratios at an early stage (epoch 10), we provide additional visualizations in Figure 8 for later stages of training. These figures illustrate the evolution of the feature space and, crucially, how the model’s representation of MEEs and other mislabeled samples changes over time.
Figure 8: Evolution of the t-SNE feature space visualization on CIFAR-10 (20% instance-dependent label noise) at (a) Epoch 40 and (b) Epoch 160. These figures complement Figure 4(a) (Epoch 10) from the main text.